# AVA-Bench: Atomic Visual Ability Benchmark for Vision Foundation Models

### v1

@@ A.1 Atomic Visual Abilities (AVAs)
    Compositional text-to-image (T2I) benchmarks. Studies on controllable generation motivate core visual primitives—number, colour, texture, object identity, spatial relations, and more—used to construct compositional prompts [36, 93]. These primitives form an initial pool of candidate abilities.
@@
    VQA question analysis. We employ GPT-4 to summarize the visual skills demanded by VQA questions in various commonly-used datasets (VQAv2 [25], RealWorldQA [94], GQA [1], etc.), thereby enriching the pool with abilities emphasized by real-world questions.
@@ A.2 Dataset Curation
Spatial Reasoning [89, 103]. We curate 15.9K image pairs from NYU-Depth V2 [74] (indoor scenes) and LVIS [29] and Objects365 [73] (open-domain scenes), all containing instance segmentation annotations. In each image, two distinct, non-overlapping objects are selected—one highlighted with a blue bounding box (reference object) and another with a red bounding box (target object). The model must identify the relative spatial position of the red box with respect to the blue box, choosing from four multiple-choice options: Left above, Left below, Right above, and Right below (Figure 11). The preprocessing steps for dataset creation are summarized below:
@@
Counting [98, 46]. We curate a total of 11.3K images from five datasets— VQAv2 [25], FSC-147 [69], CARPK [33], LVIS [29] , and CrowdHuman [48] —to evaluate object counting abilities across diverse domains, including open-domain scenes, natural objects, structured environments, and densely crowded contexts. Each image is paired with a question prompting the model to count the number of instances of a specified object category, and the model must return an integer-valued answer(Figure 12). The preprocessing steps for dataset creation are summarized below:
@@
Localization Objects365 [73] Open 27.8K 6.9K
LVIS [29] Open
iNaturalist-2021 [85] Bird, Animal
DIOR [49] Remote-Sensing
Counting VQAv2 [25] Open 9K 2.4K
FSC [69] Open
CARPK [33] Car
Crowd Surveillance [48] People
LVIS [29] Open
Fine-grained CUB-200-2011 [87] Bird 7.2K 1.8K
iNaturalist-2021 [85] Fungi,Plant, Animal
FGVC-Aircraft [59] Object
Absolute Depth NYU-Depth V2 [74] Indoor Scene 6.9K 1.8K
KITTI [23] Outdoor Scene
Relative Depth NYU-Depth V2 [74] Indoor Scene 9.2K 2.4K
KITTI [23] Outdoor Scene
OCR COCO-Text [86] Open 8.8K 2.2K
IIIT5K  [62] Open
TextVQA [26] Open
Orientation EgoOrientBench [39] Open 6.9K 1.6K
CURE-OR [77] Indoor
Object Recognition Objects365 [73] Open 27.9K 6.9K
LVIS [29] Open
iNaturalist-2021 [85] Bird, Animal
DIOR [49] Remote-Sensing
Action Recognition MiT [63] Open 12.1K 3K
Texture DTD [14] Open 10.6K 2.7K
Kylberg [45] Open
KTH-TIPS [78] Open
KTH-TIPS2 [60] Open
Spatial Reasoning Objects365 [73] Open 13K 2.9K
LVIS [29] Open
NYU-Depth V2 [74] Indoor Scene
Scene Recognition Places434 [109] Open 11.1K 2.8K
AID [95] Remote-Sensing
Emotion RAF-DB [50] Human 11.9K 5.1K
ExpW [52] Human
Color Objects365 [73] Open 11.2K 2.8K
@@
Fine-grained [66, 28, 107]. We curate a total of 9K images from five fine-grained recognition domains—Bird, Animal, Fungi, Plant, and Object—to assess species-level recognition capabilities. The dataset sources include CUB-200-2011 [87] for birds, iNat21 [85] for animals, fungi, and plants, and FGVC Aircraft [59] for objects. Each sample contains an image and a question prompting the model to identify the specific species or object type(Figure 13). Construction details are as follows:
@@
OCR [37, 55]. We curate 10.9K images from three OCR datasets—COCO-Text [86], IIIT5K [62], and TextVQA [26]—each containing word-level bounding box annotations. In each image, a red bounding box highlights the word to be transcribed. The model is prompted to recognize the textual content inside the box based on visual context(Figure 14). The preprocessing steps for dataset creation are summarized below:
@@
Localization [90, 72]. We curate 34.8K localization-focused image–question pairs sourced from Objects365 [73] and LVIS [29] (open-domain), iNaturalist-2021 [85] (birds and animals), and DIOR [49] (remote-sensing). Each image contains a single object instance, and the model is prompted to identify its location by providing the bounding box coordinates(Figure 15). The preprocessing steps for dataset creation are summarized below:
@@
    background $\text{background color}=RGB(124,120,111)$
@@
Recognition [21, 105, 83]. We curate 34.8K recognition samples from four datasets spanning diverse visual domains—Objects365 [73] and LVIS [29] (open-domain objects), iNaturalist-2021 [85] (birds and animals), and DIOR [49] (remote sensing). These samples are derived from the same images and object instances used in the localization benchmark. However, instead of asking for bounding box prediction, the recognition task requires the model to identify the object within a visually highlighted region (Figure 16).
@@
Color [88, 11]. We curate 14K images from Objects365 [73] and LVIS [29], each contributing 7K samples. This AVA focuses on assessing color perception in natural scenes. For each image, we extract a coherent color region using the following pipeline:
@@
Action [76, 88]. We curate a total of 15K image–question pairs from the Moments in Time [63] dataset, covering a wide range of human actions and activities. Each sample is derived from a short video clip annotated with a specific action label (Figure 17). Construction details are as follows:
@@
Emotion [97, 47]. We curate 16K image-question pairs from two large-scale facial expression datasets: RAF-DB [50] and ExpW [52]. These datasets consist of human portraits labeled with one of seven basic emotions: happy, sad, angry, fear, surprise, neutral, and disgust. Each image is annotated with a bounding box localizing the face of interest.
@@
Scene [15, 19]. We curate 13.9K image-question pairs from two diverse datasets: Places434 [109] (open-domain) and AID [95] (remote sensing). Each image is paired with a multiple-choice question, where the model selects the correct scene category from a pool of 30 randomly sampled options. The full set includes 464 unique scene classes spanning a wide range of environments(Figure 18). The preprocessing steps for dataset creation are summarized below:
@@
Texture [17, 22]. To assess a VFM’s ability to distinguish fine-grained visual patterns, we curate 13.2K image-question pairs from diverse surface textures using close-up images from four open-domain datasets: DTD [14] , Kylberg [45] , KTH-TIPS [78] and KTH-TIPS2 [60]. These datasets encompass a diverse array of texture types—such as striped, aluminum foil, and zigzagged—capturing subtle visual patterns that are essential for accurate texture recognition. Each image is paired with a multiple-choice question, requiring the model to select the correct texture label from a set of options(Figure 19). The preprocessing steps for dataset creation are summarized below:
@@
Orientation [99, 39]. To evaluate viewpoint understanding, we curate 10.1K image-question pairs from two specialized datasets: CURE-OR [77] and EgoOrientBench [39] . These datasets provide uncluttered images of objects captured from nine distinct orientations—front, back, left, right, top, front left, front right, back left, and back right—allowing models to learn pose-specific cues without requiring bounding boxes(Figure 20). The preprocessing steps for dataset creation are summarized below:
@@
Absolute Depth [96, 61, 104]. We curate 9K image-object pairs from NYU-Depth V2 [74] for indoor scenes and KITTI [23] for outdoor scenes. These datasets contain aligned RGB and depth information. Each image includes an object annotated with a bounding box, and the task requires estimating the absolute depth of the object in meters. This value is then matched against discretized ground-truth bins for evaluation(Figure 21).
@@
  • Image crops were extracted per object while maintaining the following aspect ratio constraint to preserve visual consistency: the width of the crop must be within the range $[\text{height},2\times\text{height}]$. This was enforced using the original image dimensions before padding or resizing.
@@
Relative Depth [96, 61, 104]. We curate 11.6K image-object pairs from NYU-Depth V2 [74] for indoor scenes and KITTI [23] for outdoor scenes, targeting the task of identifying which of two objects in an image is closer to the camera. Each image contains two distinct objects, each annotated with a bounding box. The model is asked to compare their absolute depth and choose the object that appears closer to the camera (Figure 22).
@@ B.1 Metric Details
We use CIEDE2000 [56] to calculate the color differences using the colour Python library. Specifically, we convert the predictions and ground-truths from CIE XYZ tristimulus format to CIE L*a*b* colour space with colour.XYZ_to_Lab, followed by the colour.delta_E(pred, gt, method="CIE 2000") for color differences.
@@
We use the Generalized Intersection-over-Union (GIoU [71], Equation 2):
@@
We evaluate OCR performance with Average Normalized Levenshtein Similarity (ANLS) [4]:
@@
where $\mathrm{Lev}(p,g)$ is the (Levenshtein) edit distance and |.||.|| . | denotes string length and N is the number of testing samples.
@@ C.1 Detailed Overall Results
Metric $\bm{\downarrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\downarrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\uparrow}$ $\bm{\downarrow}$
@@ C.2 Detailed Analyses for Each AVA
Fine-grained classification results are consistently robust across datasets, with the exception of SAM and MiDas, both of which are known to lack semantically rich features [18], resulting in poorer performance (see Figure 28).
@@
OCR results show consistent patterns with those reported in Section 5.2, highlighting significant underperformance by non-language-aligned VFMs, such as DINOv2 and SAM. Notably, we observe that relative performances across VFMs are stable on short texts (length $<$ 20). However, performance for CLIP and AIM sharply declines with longer text sequences (length $>$ 20), indicating potential limitations in handling extensive textual information (see Figure 30).
@@ D.1 VFM Evaluation
Existing evaluation VFM protocols generally fall into two categories. The first focuses on task-specific capabilities, typically attaching tailored heads to VFMs, followed by fine-tuning and evaluation on dedicated datasets such as ImageNet for classification [30] and COCO for detection or segmentation [79]. For example, DINOv2 [65] is evaluated on image and video classification, instance recognition, image retrieval, semantic segmentation, and depth estimation. For each task, a task-specific head is trained.
To better capture the diverse and complex perception challenges of the real world, recent studies advocate a more generic approach that leverages large language models (LLMs) as general-purpose heads, evaluating VFMs on broad Visual Question Answering (VQA) benchmarks [54, 110, 12]. For example, in addition to the traditional task-specific evaluation, AIMv2 [20] and RADIO  [70] follow the LLM-based evaluation and use a Llama-3.0(8B) [27] and a Vicuna-1.5(7B) [108] respectively, demonstrating a shift towards generalized multimodal evaluation.
@@ D.2 Atomic Visual Abilities
A recent work [7] introduced AVSBench to evaluate whether MLLMs understand basic geometric features, including angle, boundary, orthogonality, and curvature, which they refer to as atomic visual skills. However, AVSBench primarily targets geometric comprehension abilities required for geometric diagrams arising in high-school level mathematics. Moreover, AVSBench provides only test data for MLLMs without addressing potential mismatches between training and test data distributions—an issue highlighted in Section 2.2. Consequently, mispredictions in AVSBench evaluations may result from data distribution mismatches rather than genuine visual deficiencies in VFM. In contrast, AVA-Bench explicitly emphasizes atomic visual abilities essential for general visual reasoning tasks commonly encountered in real-world scenarios. By aligning training and evaluation data distributions, AVA-Bench ensures that evaluation outcomes reliably reflect genuine visual perceptual capabilities of VFMs.
Additionally, a concurrent work [92] defines a set of atomic visual capabilities analogous to ours. However, their goal fundamentally differs from ours: [92] aims to build a visual compositional tuning data recipe that builds complex capabilities from simple atomic capabilities, which can significantly reduce instruction-tuning data volume while maintaining strong performance. In contrast, AVA-Bench’s objective is to systematically evaluate VFMs against atomic visual abilities, pinpointing their exact strengths and weaknesses, and providing a comprehensive diagnostic tool to advance the development of robust vision foundation models.

### v2

@@ A.1 Atomic Visual Abilities (AVAs)
    Compositional text-to-image (T2I) benchmarks. Studies on controllable generation motivate core visual primitives—number, colour, texture, object identity, spatial relations, and more—used to construct compositional prompts huang2023t2i ; wu2024conceptmix . These primitives form an initial pool of candidate abilities.
@@
    VQA question analysis. We employ GPT-4 to summarize the visual skills demanded by VQA questions in various commonly-used datasets (VQAv2 goyal2017making , RealWorldQA RealWorldQA , GQA ainslie2023gqa , etc.), thereby enriching the pool with abilities emphasized by real-world questions.
@@ A.2 Dataset Curation
Spatial Reasoning wu2025spatial ; zhang2025call . We curate 15.9K image pairs from NYU-Depth V2 silberman2012indoor (indoor scenes) and LVIS gupta2019lvis and Objects365 shao2019objects365 (open-domain scenes), all containing instance segmentation annotations. In each image, two distinct, non-overlapping objects are selected—one highlighted with a blue bounding box (reference object) and another with a red bounding box (target object). The model must identify the relative spatial position of the red box with respect to the blue box, choosing from four multiple-choice options: Left above, Left below, Right above, and Right below (Figure 11). The preprocessing steps for dataset creation are summarized below:
@@
Counting yao2025countllm ; li2024llmcount . We curate a total of 11.3K images from five datasets— VQAv2 goyal2017making , FSC-147 ranjan2021learning , CARPK hsieh2017drone , LVIS gupta2019lvis , and CrowdHuman li2022video —to evaluate object counting abilities across diverse domains, including open-domain scenes, natural objects, structured environments, and densely crowded contexts. Each image is paired with a question prompting the model to count the number of instances of a specified object category, and the model must return an integer-valued answer(Figure 12). The preprocessing steps for dataset creation are summarized below:
@@
Localization Objects365 shao2019objects365 Open 27.8K 6.9K
LVIS gupta2019lvis Open
iNaturalist-2021 van2021benchmarking Bird, Animal
DIOR li2020object Remote-Sensing
Counting VQAv2 goyal2017making Open 9K 2.4K
FSC ranjan2021learning Open
CARPK hsieh2017drone Car
Crowd Surveillance li2022video People
LVIS gupta2019lvis Open
Fine-grained CUB-200-2011 wah2011caltech Bird 7.2K 1.8K
iNaturalist-2021 van2021benchmarking Fungi,Plant, Animal
FGVC-Aircraft maji2013fine Object
Absolute Depth NYU-Depth V2 silberman2012indoor Indoor Scene 6.9K 1.8K
KITTI geiger2013vision Outdoor Scene
Relative Depth NYU-Depth V2 silberman2012indoor Indoor Scene 9.2K 2.4K
KITTI geiger2013vision Outdoor Scene
OCR COCO-Text veit2016cocotextdatasetbenchmarktext Open 8.8K 2.2K
IIIT5K  MishraBMVC12 Open
TextVQA balanced_vqa_v2 Open
Orientation EgoOrientBench jung2024right Open 6.9K 1.6K
CURE-OR Temel2018_ICMLA Indoor
Object Recognition Objects365 shao2019objects365 Open 27.9K 6.9K
LVIS gupta2019lvis Open
iNaturalist-2021 van2021benchmarking Bird, Animal
DIOR li2020object Remote-Sensing
Action Recognition MiT monfort2019moments Open 12.1K 3K
Texture DTD cimpoi14describing Open 10.6K 2.7K
Kylberg kylberg2011kylberg Open
KTH-TIPS kth-tips_dataset Open
KTH-TIPS2 article Open
Spatial Reasoning Objects365 shao2019objects365 Open 13K 2.9K
LVIS gupta2019lvis Open
NYU-Depth V2 silberman2012indoor Indoor Scene
Scene Recognition Places434 zhou2017places Open 11.1K 2.8K
AID xia2017aid Remote-Sensing
Emotion RAF-DB li2019reliable Human 11.9K 5.1K
ExpW lian2020expression Human
Color Objects365 shao2019objects365 Open 11.2K 2.8K
@@
Fine-grained paul2023simple ; gu2025bioclip ; zhang2025finer . We curate a total of 9K images from five fine-grained recognition domains—Bird, Animal, Fungi, Plant, and Object—to assess species-level recognition capabilities. The dataset sources include CUB-200-2011 wah2011caltech for birds, iNat21 van2021benchmarking for animals, fungi, and plants, and FGVC Aircraft maji2013fine for objects. Each sample contains an image and a question prompting the model to identify the specific species or object type(Figure 13). Construction details are as follows:
@@
OCR huang2025ocr ; liu2024ocrbench . We curate 10.9K images from three OCR datasets—COCO-Text veit2016cocotextdatasetbenchmarktext , IIIT5K MishraBMVC12 , and TextVQA balanced_vqa_v2 —each containing word-level bounding box annotations. In each image, a red bounding box highlights the word to be transcribed. The model is prompted to recognize the textual content inside the box based on visual context(Figure 14). The preprocessing steps for dataset creation are summarized below:
@@
Localization wu2024visionllm ; sapkota2025object . We curate 34.8K localization-focused image–question pairs sourced from Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote-sensing). Each image contains a single object instance, and the model is prompted to identify its location by providing the bounding box coordinates(Figure 15). The preprocessing steps for dataset creation are summarized below:
@@
    $\text{background color}=RGB(124,120,111)$
@@
Recognition fu2024blink ; zhang2025revisiting ; tu2023holistic . We curate 34.8K recognition samples from four datasets spanning diverse visual domains—Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain objects), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote sensing). These samples are derived from the same images and object instances used in the localization benchmark. However, instead of asking for bounding box prediction, the recognition task requires the model to identify the object within a visually highlighted region (Figure 16).
@@
Color wang2025harnessing ; chiu2024megacoin . We curate 14K images from Objects365 shao2019objects365 and LVIS gupta2019lvis , each contributing 7K samples. This AVA focuses on assessing color perception in natural scenes. For each image, we extract a coherent color region using the following pipeline:
@@
Action szot2024grounding ; wang2025harnessing ; xie2025efficiently . We curate a total of 15K image–question pairs from the Moments in Time monfort2019moments dataset, covering a wide range of human actions and activities. Each sample is derived from a short video clip annotated with a specific action label (Figure 17). Construction details are as follows:
@@
Emotion yang2024emollm ; li2024eald . We curate 16K image-question pairs from two large-scale facial expression datasets: RAF-DB li2019reliable and ExpW lian2020expression . These datasets consist of human portraits labeled with one of seven basic emotions: happy, sad, angry, fear, surprise, neutral, and disgust. Each image is annotated with a bounding box localizing the face of interest.
@@
Scene dai2025humanvlm ; fan2024mllm . We curate 13.9K image-question pairs from two diverse datasets: Places434 zhou2017places (open-domain) and AID xia2017aid (remote sensing). Each image is paired with a multiple-choice question, where the model selects the correct scene category from a pool of 30 randomly sampled options. The full set includes 464 unique scene classes spanning a wide range of environments(Figure 18). The preprocessing steps for dataset creation are summarized below:
@@
Texture eppel2025shape ; gavrikov2024vision . To assess a VFM’s ability to distinguish fine-grained visual patterns, we curate 13.2K image-question pairs from diverse surface textures using close-up images from four open-domain datasets: DTD cimpoi14describing , Kylberg kylberg2011kylberg , KTH-TIPS kth-tips_dataset and KTH-TIPS2 article . These datasets encompass a diverse array of texture types—such as striped, aluminum foil, and zigzagged—capturing subtle visual patterns that are essential for accurate texture recognition. Each image is paired with a multiple-choice question, requiring the model to select the correct texture label from a set of options(Figure 19). The preprocessing steps for dataset creation are summarized below:
@@
Orientation yin2025multimodal ; jung2024right . To evaluate viewpoint understanding, we curate 10.1K image-question pairs from two specialized datasets: CURE-OR Temel2018_ICMLA and EgoOrientBench jung2024right . These datasets provide uncluttered images of objects captured from nine distinct orientations—front, back, left, right, top, front left, front right, back left, and back right—allowing models to learn pose-specific cues without requiring bounding boxes(Figure 20). The preprocessing steps for dataset creation are summarized below:
@@
Absolute Depth xia2024large ; mi2024hierarchical ; zhang2025vision . We curate 9K image-object pairs from NYU-Depth V2 silberman2012indoor for indoor scenes and KITTI geiger2013vision for outdoor scenes. These datasets contain aligned RGB and depth information. Each image includes an object annotated with a bounding box, and the task requires estimating the absolute depth of the object in meters. This value is then matched against discretized ground-truth bins for evaluation(Figure 21).
@@
  • Image crops were extracted per object while maintaining the following aspect ratio constraint to preserve visual consistency: the width of the crop must be within the range $\text{height},2\times\text{height}]$. This was enforced using the original image dimensions before padding or resizing.
@@
Relative Depth xia2024large ; mi2024hierarchical ; zhang2025vision . We curate 11.6K image-object pairs from NYU-Depth V2 silberman2012indoor for indoor scenes and KITTI geiger2013vision for outdoor scenes, targeting the task of identifying which of two objects in an image is closer to the camera. Each image contains two distinct objects, each annotated with a bounding box. The model is asked to compare their absolute depth and choose the object that appears closer to the camera (Figure 22).
@@ B.1 Metric Details
We use CIEDE2000 luo2001development to calculate the color differences using the colour Python library. Specifically, we convert the predictions and ground-truths from CIE XYZ tristimulus format to CIE L*a*b* colour space with colour.XYZ_to_Lab, followed by the colour.delta_E(pred, gt, method="CIE 2000") for color differences.
@@
We use the Generalized Intersection-over-Union (GIoU rezatofighi2019generalized , Equation 2):
@@
We evaluate OCR performance with Average Normalized Levenshtein Similarity (ANLS) biten2019scene :
@@
where $\mathrm{Lev}(p,g)$ is the (Levenshtein) edit distance and $|.|$ denotes string length and N is the number of testing samples.
@@ C.1 Detailed Overall Results
[…]
@@ C.2 Detailed Analyses for Each AVA
Fine-grained classification results are consistently robust across datasets, with the exception of SAM and MiDas, both of which are known to lack semantically rich features espinosa2024there ; chen2023segment , resulting in poorer performance (see Figure 28).
@@
OCR results show consistent patterns with those reported in Section 5.2, highlighting significant underperformance by non-language-aligned VFMs, such as DINOv2 and SAM. Notably, we observe that relative performances across VFMs are stable on short texts (length << 20). However, performance for CLIP and AIM sharply declines with longer text sequences (length >> 20), indicating potential limitations in handling extensive textual information (see Figure 30).
@@ D.1 VFM Evaluation
Existing evaluation VFM protocols generally fall into two categories. The first focuses on task-specific capabilities, typically attaching tailored heads to VFMs, followed by fine-tuning and evaluation on dedicated datasets such as ImageNet for classification han2022survey and COCO for detection or segmentation thisanke2023semantic . For example, DINOv2 oquab2023dinov2 is evaluated on image and video classification, instance recognition, image retrieval, semantic segmentation, and depth estimation. For each task, a task-specific head is trained.
To better capture the diverse and complex perception challenges of the real world, recent studies advocate a more generic approach that leverages large language models (LLMs) as general-purpose heads, evaluating VFMs on broad Visual Question Answering (VQA) benchmarks liu2023visual ; zhu2023minigpt ; chowdhery2023palm . For example, in addition to the traditional task-specific evaluation, AIMv2 fini2024multimodal and RADIO  ranzinger2024radio follow the LLM-based evaluation and use a Llama-3.0(8B) grattafiori2024llama and a Vicuna-1.5(7B) zheng2023judging respectively, demonstrating a shift towards generalized multimodal evaluation.
@@ D.2 Atomic Visual Abilities
A recent work chae24decomposing introduced AVSBench to evaluate whether MLLMs understand basic geometric features, including angle, boundary, orthogonality, and curvature, which they refer to as atomic visual skills. However, AVSBench primarily targets geometric comprehension abilities required for geometric diagrams arising in high-school level mathematics. Moreover, AVSBench provides only test data for MLLMs without addressing potential mismatches between training and test data distributions—an issue highlighted in Section 2.2. Consequently, mispredictions in AVSBench evaluations may result from data distribution mismatches rather than genuine visual deficiencies in VFM. In contrast, AVA-Bench explicitly emphasizes atomic visual abilities essential for general visual reasoning tasks commonly encountered in real-world scenarios. By aligning training and evaluation data distributions, AVA-Bench ensures that evaluation outcomes reliably reflect genuine visual perceptual capabilities of VFMs.
Additionally, a concurrent work wu2025compact defines a set of atomic visual capabilities analogous to ours. However, their goal fundamentally differs from ours: wu2025compact aims to build a visual compositional tuning data recipe that builds complex capabilities from simple atomic capabilities, which can significantly reduce instruction-tuning data volume while maintaining strong performance. In contrast, AVA-Bench’s objective is to systematically evaluate VFMs against atomic visual abilities, pinpointing their exact strengths and weaknesses, and providing a comprehensive diagnostic tool to advance the continual development mai2022online ; lomonaco2022cvpr ; mai2021supervised ; shim2021online ; mai2020batch of robust vision foundation models.